# New supersymmetric $AdS_{6}$ black holes from matter-coupled $F(4)$ gauged supergravity

### v2

@@ Abstract
We study supersymmetric $AdS_{6}$ black holes from matter-coupled $N=(1,1)$ gauged supergravity coupled to three vector multiplets and $SO(3)\times SO(3)$ gauge group. This gauged supergravity admits two supersymmetric $AdS_{6}$ vacua preserving all supersymmetries with $SO(3)\times SO(3)$ and $SO(3)_{\textrm{diag}}$ symmetries. By considering a truncation to $SO(2)_{\textrm{diag}}\subset SO(3)_{\textrm{diag}}$ invariant sector, we find a number of new supersymmetric $AdS_{2}\times\mathcal{M}_{4}$ solutions by performing a topological twist along $\mathcal{M}_{4}$. For $\mathcal{M}_{4}$ being a product of two Riemann surfaces $\Sigma\times\widetilde{\Sigma}$ and a Kahler four-cycle, the twist is implemented by $SO(2)_{\textrm{diag}}$ gauge field while, for $\mathcal{M}_{4}$ given by a Clayley four-cycle, the twist is performed by turning on $SO(3)_{\textrm{diag}}$ gauge fields. We give numerical black hole solutions interpolating between $AdS_{2}\times\mathcal{M}_{4}$ near horizon geometries and asymptotically locally $AdS_{6}$ vacua. Among these solutions, there are solutions interpolating between both of the supersymmetric $AdS_{6}$ vacua and near horizon geometries. The solutions can also be interpreted as holographic RG flows from five-dimensional SCFTs to superconformal quantum mechanics.
@@ 1 Introduction
In six dimensions, only the half-maximal non-chiral $N=(1,1)$ or $F(4)$ gauged supergravity constructed in [19], see also [20, 21] for the matter-coupled $F(4)$ gauged supergravity, admits supersymmetric $AdS_{6}$ vacua, see [22] for general conditions on the existence of supersymmetric $AdS_{6}$ vacua. The corresponding $AdS_{6}$ black hole solutions with the near horizon geometry of the form $AdS_{2}\times\mathcal{M}_{4}$ from pure $F(4)$ gauged supergravity has been considered in [23] and more recently in [24] and [25]. More general solutions in the matter-coupled $F(4)$ gauged supergravity have appeared later in [26] and [27]. In [26], the $F(4)$ gauged supergravity is coupled to one vector multiplet leading to $SO(3)\times U(1)$ gauge group with all the fields being uncharged under the $U(1)$ factor. A more general $F(4)$ gauged supergravity coupled to three vector multiplets has been considered in [27]. The resulting gauged supergravity has $SO(3)\times SO(3)$ gauge group. However, the black hole solutions with $AdS_{2}\times H^{2}\times H^{2}$ and $AdS_{2}\times\mathcal{M}_{4}$ for $\mathcal{M}_{4}$ being a Kahler four-cycle given in [27] have been found by considering a truncation to $SO(2)\times SO(2)$ invariant sector. Within this sector, the effect of the second $SO(3)$ factor in the gauge group is invisible. In particular, the gauge coupling constant for this factor does not appear at all in the BPS equations.
@@
In this work, we look for more interesting supersymmetric $AdS_{6}$ black hole solutions within this matter-coupled $F(4)$ gauged supergravity with three vector multiplets and $SO(3)\times SO(3)$ gauge group. This gauged supergravity has been first studied in [28], and further in [29, 30], in which two $N=(1,1)$ supersymmetric $AdS_{6}$ vacua with $SO(3)\times SO(3)$ and $SO(3)_{\textrm{diag}}$ symmetries have been found. By further reduce the residual symmetry to $SO(2)_{\textrm{diag}}$, we will look for black holes with hear horizon geometries of the form $AdS_{2}\times\mathcal{M}_{4}$ for $\mathcal{M}_{4}$ being a product of two Riemann surfaces or a Kahler four-cycle. We will see that, for negatively curved $\mathcal{M}_{4}$ spaces, there exist a number of supersymmetric black hole solutions in asymptotically $AdS_{6}$ spaces described by the two aforementioned $AdS_{6}$ vacua. In addition, we will also consider the black hole solutions with $\mathcal{M}_{4}$ given by a Cayley four-cycle by performing a topological twist using $SO(3)_{\textrm{diag}}$ gauge fields. Although this has already been studied in [27], we indeed find new supersymmetric $AdS_{6}$ black hole solutions interpolating between the near horizon geometry and the two supersymmetric $AdS_{6}$ vacua.
@@ 2.1 General structure of matter-coupled $F(4)$ gauged supergravity
This gauged supergravity has been originally studied in [28], and two supersymmetric $N=(1,1)$ $AdS_{6}$ vacua with $SO(3)\times SO(3)$ and $SO(3)_{\textrm{diag}}$ symmetries have been identified. For convenience, we will also give these two vacua here. With the $SO(3)_{\textrm{diag}}$ generated by $J^{rs}+\tilde{J}^{rs}$, the only one singlet scalar from $SO(4,3)/SO(4)\times SO(3)$ coset corresponds to the non-compact generator $Y_{11}+Y_{22}+Y_{33}$. The coset representative can be written as
@@ 3 Supersymmetric $AdS_{6}$ black holes with $\Sigma\times\widetilde{\Sigma}$ horizons
with hatted indices being flat indices. Throughout the paper, we will use ${}^{\prime}$ to denote $r$-derivatives except for $F^{\prime}_{\kappa}(\theta)=\frac{dF_{\kappa}(\theta)}{d\theta}$ and $\tilde{F}^{\prime}_{\tilde{\kappa}}(\tilde{\theta})=\frac{d\tilde{F}_{\tilde{\kappa}}(\tilde{\theta})}{d\tilde{\theta}}$. We also note some useful relations
@@ 3.2 Numerical black hole solutions
Figure 2: Supersymmetric $AdS_{6}$ black holes interpolating between $AdS_{6}$ vacua $i$ and $ii$ and the near horizon geometry $AdS_{2}\times H^{2}\times H^{2}$ (critical point II) for $g_{2}$=4.
@@ 4.1 Black holes with Kahler four-cycle horizon
Examples of black hole solutions with the near horizon geometry given by critical point IV are shown in figure 4 for $\phi_{3}=0$ and $g_{2}=4$. As in the previous section, there exist solutions interpolating between the two supersymmetric $AdS_{6}$ vacua (critical point $i$ and $ii$) and $AdS_{2}\times\mathcal{M}^{-}_{4}$ geometry IV. In this case, the entropy of the black hole is given by
@@
Figure 4: Supersymmetric $AdS_{6}$ black holes interpolating between $AdS_{6}$ vacua $i$ and $ii$ and the near horizon geometry $AdS_{2}\times\mathcal{M}^{-}_{4}$ (critical point IV) for $g_{2}=4$.
@@ 4.2 Black holes with Cayley four-cycle horizon
As in all the other cases, the solution only exists for $\kappa=-1$. Examples of numerical solutions for $m=\frac{1}{2}$ and $g_{2}=4$ are given in figure 5. The entropy of the black hole is given by
@@
Figure 5: Supersymmetric $AdS_{6}$ black holes interpolating between $AdS_{6}$ vacua $i$ and $ii$ and the near horizon geometry $AdS_{2}\times\mathcal{M}_{4}^{-}$ (critical point V) for $m=\frac{1}{2}$ and $g_{2}=4$.
@@ 5 Conclusions
We have constructed new supersymmetric $AdS_{6}$ black hole solutions from $F(4)$ gauged supergravity in six dimensions coupled to three vector multiplets with $SO(3)\times SO(3)$ gauge group. By considering a truncation to $SO(2)_{\textrm{diag}}$ singlet sector and performing a topological twist by the $SO(2)_{\textrm{diag}}$ gauge field, we find a number of black hole solutions with $AdS_{2}\times H^{2}\times H^{2}$ and $AdS_{2}\times\mathcal{M}^{-}_{4}$ near horizon geometries with $\mathcal{M}_{4}^{-}$ being a negatively curved Kahler four-cycle. On the other hand, by performing a twist by $SO(3)_{\textrm{diag}}$ gauge fields, we have found a new black hole solution with $AdS_{2}\times\mathcal{M}_{4}^{-}$ near horizon geometry for $\mathcal{M}^{-}_{4}$ being a negatively curved Cayley four-cycle.
@@ Appendix A Bosonic field equations of matter-coupled $F(4)$ gauged supergravity
$\left.+\frac{3}{32}e^{4\sigma}H_{\rho\sigma\lambda}H^{\rho\sigma\lambda}+2\partial_{\rho}\sigma\partial^{\rho}\sigma+\frac{1}{2}P^{I\alpha}_{\rho}P^{\rho}_{I\alpha}-2V\right]$ $=$ 00 0$\qquad$ (112)

### v3

@@ Abstract
We study supersymmetric $AdS_{6}$ black holes from matter-coupled $N=(1,1)$ gauged supergravity coupled to three vector multiplets and $SO(3)\times SO(3)$ gauge group. This gauged supergravity admits two supersymmetric $AdS_{6}$ vacua preserving all supersymmetries with $SO(3)\times SO(3)$ and $SO(3)_{\textrm{diag}}$ symmetries. By considering a truncation to $SO(2)_{\textrm{diag}}\subset SO(3)_{\textrm{diag}}$ invariant sector, we find a number of new supersymmetric $AdS_{2}\times\mathcal{M}_{4}$ solutions by performing a topological twist along $\mathcal{M}_{4}$. For $\mathcal{M}_{4}$ being a product of two Riemann surfaces $\Sigma\times\widetilde{\Sigma}$ and a Kahler four-cycle, the twist is implemented by $SO(2)_{\textrm{diag}}$ gauge field while, for $\mathcal{M}_{4}$ given by a Cayley four-cycle, the twist is performed by turning on $SO(3)_{\textrm{diag}}$ gauge fields. We give numerical black hole solutions interpolating between $AdS_{2}\times\mathcal{M}_{4}$ near horizon geometries and asymptotically locally $AdS_{6}$ vacua. Among these solutions, there are solutions interpolating between both of the supersymmetric $AdS_{6}$ vacua and near horizon geometries. The solutions can also be interpreted as holographic RG flows from five-dimensional SCFTs to superconformal quantum mechanics.
@@ 1 Introduction
In six dimensions, only the half-maximal non-chiral $N=(1,1)$ or $F(4)$ gauged supergravity constructed in [19] admits supersymmetric $AdS_{6}$ vacua. The matter-coupled $F(4)$ gauged supergravity has been constructed in [20, 21], and general conditions on the existence of supersymmetric $AdS_{6}$ vacua have been given in [22]. The corresponding $AdS_{6}$ black hole solutions with the near horizon geometry of the form $AdS_{2}\times\mathcal{M}_{4}$ from pure $F(4)$ gauged supergravity has been considered in [23] and more recently in [24] and [25]. More general solutions in the matter-coupled $F(4)$ gauged supergravity have appeared later in [26] and [27]. In [26], the $F(4)$ gauged supergravity is coupled to one vector multiplet leading to $SO(3)\times U(1)$ gauge group with all the fields being uncharged under the $U(1)$ factor. A more general $F(4)$ gauged supergravity coupled to three vector multiplets has been considered in [27]. The resulting gauged supergravity has $SO(3)\times SO(3)$ gauge group. However, the black hole solutions with $AdS_{2}\times H^{2}\times H^{2}$ and $AdS_{2}\times\mathcal{M}_{4}$ for $\mathcal{M}_{4}$ being a Kahler four-cycle given in [27] have been found by considering a truncation to $SO(2)\times SO(2)$ invariant sector. Within this sector, the effect of the second $SO(3)$ factor in the gauge group is invisible. In particular, the gauge coupling constant for this factor does not appear at all in the BPS equations.
@@
In this work, we look for more interesting supersymmetric $AdS_{6}$ black hole solutions within this matter-coupled $F(4)$ gauged supergravity with three vector multiplets and $SO(3)\times SO(3)$ gauge group. This gauged supergravity has been first studied in [28], and further in [29], in which two $N=(1,1)$ supersymmetric $AdS_{6}$ vacua with $SO(3)\times SO(3)$ and $SO(3)_{\textrm{diag}}$ symmetries have been found together with holographic RG flows between $AdS_{6}$ vacua and RG flows to non-conformal phases of the dual five-dimensional SCFTs. In particular, supersymmetric $AdS_{4}\times\Sigma$ with $\Sigma=S^{2},H^{2}$ and $AdS_{3}\times\mathcal{M}_{3}$ with $\mathcal{M}_{3}=S^{3},H^{3}$ solutions have been studied in [30]. These solutions correspond to near horizon geometries of black strings and black two-branes in asymptotically $AdS_{6}$ spaces.
In contrast to the previous study of $SO(2)\times SO(2)$ invariant sector, we will consider a smaller residual symmetry $SO(2)_{\textrm{diag}}\subset SO(2)\times SO(2)$ and look for black holes with near horizon geometries of the form $AdS_{2}\times\mathcal{M}_{4}$ for $\mathcal{M}_{4}$ being a product of two Riemann surfaces or a Kahler four-cycle. We will see that, for negatively curved $\mathcal{M}_{4}$ spaces, there exist a number of supersymmetric black hole solutions in asymptotically $AdS_{6}$ spaces described by the two aforementioned $AdS_{6}$ vacua. In addition, we will also consider the black hole solutions with $\mathcal{M}_{4}$ given by a Cayley four-cycle by performing a topological twist using $SO(3)_{\textrm{diag}}$ gauge fields. Although this has already been studied in [27], we indeed find new supersymmetric $AdS_{6}$ black hole solutions interpolating between the near horizon geometry and the two supersymmetric $AdS_{6}$ vacua.
@@ 2.1 General structure of matter-coupled $F(4)$ gauged supergravity
This gauged supergravity has been originally studied in [28], and two supersymmetric $N=(1,1)$ $AdS_{6}$ vacua with $SO(3)\times SO(3)$ and $SO(3)_{\textrm{diag}}$ symmetries have been identified. For convenience, we will also present these two vacua here. With the $SO(3)_{\textrm{diag}}$ generated by $J^{rs}+\tilde{J}^{rs}$, the only one singlet scalar from $SO(4,3)/SO(4)\times SO(3)$ coset corresponds to the non-compact generator $Y_{11}+Y_{22}+Y_{33}$. The coset representative can be written as
@@ 3 Supersymmetric $AdS_{6}$ black holes with $\Sigma\times\widetilde{\Sigma}$ horizons
with hatted indices being flat indices. Throughout the paper, we will use to denote $r$-derivatives except for $F^{\prime}_{\kappa}(\theta)=\frac{dF_{\kappa}(\theta)}{d\theta}$ and $\tilde{F}^{\prime}_{\tilde{\kappa}}(\tilde{\theta})=\frac{d\tilde{F}_{\tilde{\kappa}}(\tilde{\theta})}{d\tilde{\theta}}$. We also note some useful relations
@@ 3.2 Numerical black hole solutions
Figure 2: Supersymmetric $AdS_{6}$ black holes interpolating between $AdS_{6}$ vacua $i$ and $ii$ and the near horizon geometry $AdS_{2}\times H^{2}\times H^{2}$ (critical point II) for $g_{2}$=4. The red and green curves represent respectively solutions that flows directly from $AdS_{6}$ critical points $i$ and $ii$ to $AdS_{2}\times H^{2}\times H^{2}$ vacuum II. Solutions interpolating between $AdS_{6}$ critical point $i$ and $AdS_{2}\times H^{2}\times H^{2}$ vacuum II that flow very close to $AdS_{6}$ vacuum $ii$ are given by blue, purple, and cyan curves. Note also that in subfigures (d), (e), and (f), solutions represented by blue and purple lines are very close to each other rendering the blue line invisible.
@@ 4.1 Black holes with Kahler four-cycle horizon
Examples of black hole solutions with the near horizon geometry given by critical point IV are shown in figure 4 for $\phi_{3}=0$ and $g_{2}=4$. As in the previous section, there are solutions flowing directly from $AdS_{6}$ critical point $i$ to $AdS_{2}\times\mathcal{M}^{-}_{4}$ fixed point IV as shown by the purple line. In addition, there exist solutions interpolating between the two supersymmetric $AdS_{6}$ vacua (critical point $i$ and $ii$) and $AdS_{2}\times\mathcal{M}^{-}_{4}$ geometry IV given by red, green, and blue lines. In this case, the entropy of the black hole is given by
@@
Figure 4: Supersymmetric $AdS_{6}$ black holes interpolating between $AdS_{6}$ vacua $i$ and $ii$ and the near horizon geometry $AdS_{2}\times\mathcal{M}^{-}_{4}$ (critical point IV) for $g_{2}=4$. The purple line represents a solution flowing directly from $AdS_{6}$ critical point $i$ to $AdS_{2}\times\mathcal{M}^{-}_{4}$ fixed point IV. The red, green, and blue lines correspond to solutions interpolating between two supersymmetric $AdS_{6}$ vacua (critical point $i$ and $ii$) and $AdS_{2}\times\mathcal{M}^{-}_{4}$ geometry IV.
@@ 4.2 Black holes with Cayley four-cycle horizon
As in all the other cases, the solution only exists for $\kappa=-1$. Examples of numerical solutions for $m=\frac{1}{2}$ and $g_{2}=4$ are given in figure 5. Similar to the previous cases, there are solutions interpolating between $AdS_{6}$ critical point $i$ and the $AdS_{2}\times\mathcal{M}_{4}^{-}$ critical point V represented by the blue curve. The orange, green, red, and purple lines correspond to solutions that flow from $AdS_{6}$ critical point $i$ and approach $AdS_{6}$ critical point $ii$ before flowing to the $AdS_{2}\times\mathcal{M}_{4}^{-}$ critical point V.
The entropy of the black hole is given by
@@
Figure 5: Supersymmetric $AdS_{6}$ black holes interpolating between $AdS_{6}$ vacua $i$ and $ii$ and the near horizon geometry $AdS_{2}\times\mathcal{M}_{4}^{-}$ (critical point V) for $m=\frac{1}{2}$ and $g_{2}=4$. A solution interpolating between $AdS_{6}$ critical point $i$ and the $AdS_{2}\times\mathcal{M}_{4}^{-}$ critical point V corresponds to the blue curve while solutions that approach $AdS_{6}$ critical point $ii$ before flowing to the $AdS_{2}\times\mathcal{M}_{4}^{-}$ critical point V are shown by orange, green, red, and purple lines.
@@ 5 Conclusions
We have constructed new supersymmetric $AdS_{6}$ black hole solutions from $F(4)$ gauged supergravity in six dimensions coupled to three vector multiplets with $SO(3)\times SO(3)$ gauge group. By considering a truncation to $SO(2)_{\textrm{diag}}$ singlet sector and performing a topological twist by the $SO(2)_{\textrm{diag}}$ gauge field, we find a number of black hole solutions with $AdS_{2}\times H^{2}\times H^{2}$ and $AdS_{2}\times\mathcal{M}^{-}_{4}$ near horizon geometries with $\mathcal{M}_{4}^{-}$ being a negatively curved Kahler four-cycle. On the other hand, by performing a twist by $SO(3)_{\textrm{diag}}$ gauge fields, we have found a new black hole solution with $AdS_{2}\times\mathcal{M}_{4}^{-}$ near horizon geometry for $\mathcal{M}^{-}_{4}$ being a negatively curved Cayley four-cycle. All the solutions identified in this paper are collectively shown in table 1.
Solution Near horizon geometry Numerical black hole solution
Critical point I $AdS_{2}\times H^{2}\times H^{2}$ (52) Figure 1
Critical point II $AdS_{2}\times H^{2}\times H^{2}$ (54) Figure 2
Critical point III $AdS_{2}\times\mathcal{M}^{-}_{\textrm{K}4}$ (81) Figure 3
Critical point IV $AdS_{2}\times\mathcal{M}^{-}_{\textrm{K}4}$ (83) Figure 4
Critical point V $AdS_{2}\times\mathcal{M}^{-}_{\textrm{C}4}$ (106) Figure 5
Table 1: Near horizon $AdS_{2}\times\mathcal{M}_{4}$ geometries and numerical solutions for supersymmetric $AdS_{6}$ black holes found in this paper. The notations $\mathcal{M}^{-}_{\textrm{K}4}$ and $\mathcal{M}^{-}_{\textrm{C}4}$ correspond to $\mathcal{M}_{4}$ being Kahler 4-cycle and Cayley four-cycle, respectively.
@@ Appendix A Bosonic field equations of matter-coupled $F(4)$ gauged supergravity
$\left.+\frac{3}{32}e^{4\sigma}H_{\rho\sigma\lambda}H^{\rho\sigma\lambda}+2\partial_{\rho}\sigma\partial^{\rho}\sigma+\frac{1}{2}P^{I\alpha}_{\rho}P^{\rho}_{I\alpha}+2V\right]$ $=$ 00 0$\qquad$ (113)
@@
Data Availability Statement: No data associated in the manuscript